OpenStudy (anonymous):

Suppose an isosceles triangle ABC has A pi/4 = and b = 4 = c. What is the length of a^2? a. 4^2(2-sqrt2) b. 4^2 sqrt2 c. 4^2(2+sqrt2) d. 4^2(sqrt2-2)

OpenStudy (hoblos):

do you know the law of cosine ?

OpenStudy (anonymous):

no

OpenStudy (hoblos):

\[a ^{2} = b^{2} + c^{2} - 2bc \cos \alpha \]

OpenStudy (hoblos):

so you have to apply this formula to calculate a^2

OpenStudy (anonymous):

when it says and b=4=c does that mean c also equals 4?

OpenStudy (hoblos):

yes

OpenStudy (anonymous):

so it would be \[a^2=4^2+4^2-2(4)(4)\cos a\]

OpenStudy (anonymous):

is the last a my (pi/4)?

OpenStudy (hoblos):

yes, and it is cos(pi/4)

OpenStudy (hoblos):

yes it is

OpenStudy (anonymous):

i got 9.37

OpenStudy (hoblos):

i don't think you have to calculate the exact value, your answer must be on of the choices given

OpenStudy (anonymous):

so its a

OpenStudy (hoblos):

correct